# Persistent Homology for High-dimensional Data Based on Spectral Methods

### v2

@@ Appendix D Effective resistance integrates diffusion distances
This is an application of the geometric series. Since $G$ is not bipartite, the largest eigenvalue $\mu_{n}$ of its Laplacian $L^{\text{sym}}$ is smaller than $2$ [16, Lemma 1.7]. Recall that by Eq. (4) the diffusion distance for $t$ diffusion steps is given by
@@
Corollary D.2.
The sum of the squared diffusion distances over all time points $t/2$ for $t=2,3,\dots$ equals the corrected effective resistance. Formally, if $G$ is a connected graph and not bipartite, we have
@@
For $x\in(-1,1)$, starting a geometric series at the $m$-th term yields $\sum_{t=m}^{\infty}x^{t}=\frac{x^{m}}{1-x}$. By Proposition C.2, we have
@@
By the same argument as in the proof of Proposition D.1, this shows the first part of the Corollary.
The statement for diffusion distance after $t=0$ steps is an easy computation from the original definition (2):
@@
and the expression of the naive effective resistance as full geometric series of squared diffusion distances in Proposition D.1. ∎
Appendix E Diffusion pseudotime
@@
The difference is the $v_{1},\dots,v_{n}$’s. In the original publication [31], they were the normalized eigenvectors of the random walker graph Laplacian $D^{-\frac{1}{2}}L^{\text{sym}}D^{\frac{1}{2}}$. These are given by $v_{l}=D^{-\frac{1}{2}}u_{l}/\|D^{-\frac{1}{2}}u_{l}\|$. Wolf et al. [75] introduced a version using the eigenvectors of the symmetric graph Laplacian $L^{\text{sym}}$, so that $v_{l}=u_{l}$. Closest to the corrected resistance distance is the case where $v_{l}=D^{-\frac{1}{2}}u_{l}$ are non-normalized. We refer to the three versions as “rw”, “sym” and “symd”.
The only difference between the “symd" version of DPT and corrected effective resistance is that the former decays eigenvalues with $(1-\mu_{l})/\mu_{l}$, while the latter decays it with $(1-\mu_{l})/\sqrt{\mu_{l}}$, so that DPT decays large eigenvalues more strongly than corrected effective resistance (Figure 4).
@@
E.1 Performance of diffusion pseudotime and potential distance
We additionally compare the spectral methods of our benchmark to the three versions of DPT and potential distance.
For DPT, we used all three versions and used $k\in\{15,100\}$ nearest neighbors.
@@ E.1 Performance of diffusion pseudotime and potential distance
Overall, all spectral methods performed very similarly on the 1D datasets circle, linked circles, and eyeglasses (Figures S3, S4), as well as on the single-cell datasets (Figure S5).
@@
Figure S5: Comparison of all spectral methods on the single cell datasets. They achieve comparative performance as effective resistance and diffusion distances.
Appendix F Details on the distances used in our benchmark
@@
$\theta=\begin{cases}(\text{dtm}_{i}+\text{dtm}_{j}+d_{ij})/2&\text{ if }\xi=1\\ \sqrt{((\text{dtm}_{i}+\text{dtm}_{j})^{2}+d_{ij}^{2})\cdot((\text{dtm}_{i}-\text{dtm}_{j})^{2}+d_{ij}^{2})}/(2d_{ij})&\text{ if }\xi=2\\ \max(\text{dtm}_{i},\text{dtm}_{j},d_{ij}/2)&\text{ if }\xi=\infty.\end{cases}$ (61)
@@
We computed the effective resistance on the $\text{s}k\text{NN}$ graph. Following the analogy with resistances in an electric circuit, if the $\text{s}k\text{NN}$ graph is disconnected, we computed the effective resistance separately in each connected component and set resistances between components to $\infty$. The uncorrected resistances were computed via the pseudoinverse of the unnormalized graph Laplacian
@@
where $l_{ij}^{\dagger}$ is the $ij$-th entry of the pseudoinverse of the unnormalized $\text{s}k\text{NN}$ graph Laplacian. The pseudoinverse inverts all non-zero eigenvalues. This is where the scaling by $\frac{1}{\sqrt{\mu_{l}}}$ comes from in Eq. 5.
@@
For the weighted version of effective resistance, each edge in the $\text{s}k\text{NN}$ graph was weighted by the inverse of the Euclidean distance. We experimented with the weighted and unweighted versions, but only reported the unweighted version in the paper (as the difference was always minor). We also experimented with the unweighted and uncorrected version and saw that correcting is crucial for high noise levels (Figure S16). We used $k~{}\in\{15,100\}$.
Both forms of the effective resistance can be written as squared distances between certain embedding points (5), (6). Nevertheless, the uncorrected effective resistance is a proper metric [27, Corollary 2.4.]. The corrected version is not a proper metric in general (Proposition C.3).
@@
It is clear from the definition of the diffusion distance (68) that it satisfies the triangle inequality, but it can fail to be positive on distinct points (Proposition C.3).
@@
Diffusion pseudotime [31] also integrates the different time scales of the diffusion distance, but on the level of the transition matrices, rather than on the level of the distances themselves (Proposition D.2). There are multiple variants of diffusion pseudotime. They are all computed as $d^{\textnormal{dpt}}_{ij}=\|e_{i}^{\textnormal{dpt}}-e_{j}^{\textnormal{dpt}}\|$ for
@@
Appendix G Datasets
G.1 Synthetic datasets
@@
G.2 Single-cell datasets
@@
Appendix H Hyperparameter selection
@@
The hyperparameters for Figures S6 and S7 are the same as those used in Figure 5. The hyperparameter for Figure S12 are given in Table S3. All other supplementary figures either do not depend on hyperparameters or detail them directly in the figure or the caption.
@@
Appendix I Hyperparameter sensitivity
While all distances other than the Euclidean distance have hyperparameters and we show results for the best hyperparameter setting in most figures, hyperparameter selection does not pose a serious problem for diffusion distance and effective resistance. We only tuned hyperparameters very mildly for these methods (4 settings for diffusion distances, only 2 for effective resistance, as opposed to 24 for the competing method DTM). Moreover, we conduced a more fine-grained sensitivity analysis and found the performance of diffusion distances and effective resistances to be robust to the exact value of their hyperparameters. For this sensitivity analysis we computed the area under the noise-level / detection score curves for the circle, the interlinked circles and the eyeglasses dataset in $50$ ambient dimensions with $k=4,15,30,45,60,75,90,105,120,135,150$ nearest neighbors and $t=2,4,8,16,32,64,128,256$ diffusion steps. For all but the most extreme hyperparameter values both effective resistance and diffusion distances outperformed the Euclidean distance significantly (Figure S11). We observed a trade-off between the optimal number of neighbors $k$ and diffusion steps $t$ for the diffusion distance, since both control how fast the diffusion can progress across the dataset.
@@ Appendix I Hyperparameter sensitivity
Appendix J Implementation details
We computed persistent homology using the ripser [4] project’s representative-cycles branch at commit 140670f to compute persistent homologies and representative cycles. We used coefficients in $\mathbb{Z}/2\mathbb{Z}$. To compute $k$NN graphs, we used the PyKeops package [12]. The rest of our implementation is in Python. Our code is available at https://github.com/berenslab/eff-ph/tree/arxiv-v2.
@@
The run time of persistent homology vastly dominated the time taken by the distance computation. The persistent homology run time depended most strongly on the sample size $n$, the dataset, and on the highest dimensionality of holes. The difference between distances was usually small. However, we observed that there were some outliers, depending on the noise level and the random seed, that had much longer run time. Overall, we found that methods that produce many pairwise distances of the same value (e.g., because of infinite distance in the graph affinities or maximum operations like for DTM with $p=\infty,\xi=\infty$) often had a much longer run time than other settings. We presume this was because equal distances lead to many simplices being added to the complex at the same time. We give exemplary run times in Table S4.
@@
Appendix K Effect of outliers
@@
In low dimensionality, Euclidean distance suffered in the low Gaussian noise setting already when only $50$ outliers were added. Adding $100$ and $200$ severely lowered the detection score for Euclidean distance across the entire Gaussian noise range. Fermat distances suffered from high random seed variability when adding outliers in low ambient dimension. Diffusion distance and effective resistance were much more outlier-resistant than the Euclidean distance and changed their performance noticeably only after $200$ outliers were added. Even then they performed better than the Euclidean distance without outliers. DTM excelled in this setting, being completely insensitive to outliers and achieving top score for all noise levels (Figure S12 a – d).
The volume of the bounding box in $d=50$ ambient dimensions is much larger and thus the same number outlier points lie much more sparsely. In particular, it is much less likely that an outlier happens to fall into the middle of the circle. As a result, even Euclidean and Fermat distance were very outlier-robust in $d=50$ ambient dimensions (Figure S12 e – g). Similarly, DTM’s performance did not change at all in the face of outliers. However, all three methods suffered strongly from the high-dimensional Gaussian noise.
As diffusion distance and effective resistance rely on an unweighted $k$NN graph in our implementation, they are somewhat more susceptible to outliers. Diffusion distance’s performance decreased steadily with the number of outliers in high ambient dimension. When outliers are present, it performed worse than the Euclidean distance in the low Gaussian noise setting, but still much better in the high Gaussian noise setting, even for $200$ outliers. Effective resistance performed best overall, deteriorating only slightly in the low Gaussian noise setting when outliers are added. Both spectral methods clearly outperformed the others in the high-Gaussian noise regime even in the face of numerous outliers.
@@
Figure S12: Loop detection performance of various methods on the noisy circle in the presence of outliers in low- and high-dimensional ambient space. Outliers are sampled uniformly from an axis-aligned cube around the data. a – d. In low ambient dimension ($d=2$) adding outliers hurt the performance of the Euclidean and Fermat distance, but barely that of the spectral methods and not at all DTM’s excellent performance. e – g. In high ambient dimension ($d=50$) outliers did not further decrease the weak performance of non-spectral methods. Diffusion distances were somewhat outlier sensitive, but could still detect the loop structure in the high Gaussian noise setting well. Effective resistance performed best overall and was not very outlier-sensitive.
Appendix L UMAP with higher embedding dimension
@@
Following its original publication [45], UMAP is typically used to embed data into two dimensions. This is an obvious issue for datasets which are not embeddable into two dimensions, such as the sphere or the torus. Therefore, we also experimented with the less common setting of higher embedding dimensions ($3$, $5$, $10$) for UMAP. Save for the successful detection of the sphere’s void with at least three embedding dimensions, we observed few consistent effects of the embedding dimension either on the toy (Figure S13) or on the real data (Figure S15). This is in contrast to Laplacian Eigenmaps for which a higher embedding dimensions has usually a strong and on most datasets negative impact (Figures S19S29). Nevertheless, on the linked circles dataset, which is not embeddable into two dimensions, we saw a small improvement for UMAP when going beyond two embedding dimensions for the low noise setting.
@@
Independent of the embedding dimension, UMAP struggled in the low-noise setting on the eyeglasses dataset for $k=15$. Moreover, we observed very poor performance both for loop and void detection on the torus. We believe UMAP’s tendency to over-fragment [20] and UMAP’s use of a heavy-tailed kernel to be the causes of this. We visualized three-dimensional UMAP embeddings of the noiseless eyeglasses, sphere, and torus for $k=15$ in Figure S14. UMAP left a gap in the embedding of the eyeglasses dataset, so that the true loop only had the second highest persistence. Short-cutting at the bottleneck yielded the most persistent loop. For the torus, the surface of the embedding was very fragmented, preventing the detection of any void. While the main loop of the torus was detected well, the second most persistent detected loop was already in the noise cloud as the fragmented surface of the embedding allowed for many fairly persistent loops. In a similar way, the surface of the sphere’s embedding got fragmented, leading to many loops. Our metric does not penalize this, because the detection score for $m=1$ loop is low as there is nearly no gap in persistence between the most and second-most persistent loop. Higher levels of noise or of $k$ seemed to overcome UMAP’s over-fragmentation tendency for the eyeglasses dataset. This might be due to high-dimensional noise matching UMAP’s heavy-tailed kernel better. However, it did not resolve the over-fragmentation for the torus.
While higher embedding dimensions impacted the performance of Laplacian Eigenmaps, varying the embedding dimension for UMAP only had a much smaller effect on toy data and no consistent effect on the real data (Figure S15).
@@
UMAP scales linearly with the embedding dimension, but typical $t$-SNE implementations scale exponentially with the embedding dimension. In fact, the one we used, openTSNE [55], does not even implement embedding dimensions higher than two, which is why we only explored higher embedding dimensions for UMAP.
@@
Figure S15: Detection scores for UMAP with different embedding dimensions on the single-cell datasets with some other methods for reference. Changing the embedding dimension did not have a consistent effect on the detection scores, while higher embedding dimension ususally hurt for Laplacian Eigenmaps.
Appendix M Additional figures
Figure S16: Loop detection score on noisy $S^{1}\subset\mathbb{R}^{50}$ for various versions of effective resistance. There was little difference between using the weighted $k$NN graph, unweighted $k$NN graph, and using the square root of effective resistance based on the unweighted $k$NN graph. The latter got filtered out for high noise levels. Using $k=100$ instead of $k=15$ helped only marginally on this dataset. The uncorrected (naive) version of effective resistance collapsed at very small noise levels.
Figure S17: Illustration for the random seed variability of effective resistance with $k=15$ on the noisy eyeglasses dataset in $\mathbb{R}^{50}$ with $\sigma=0.075$. This refers to panel b of Figure 6. a. One-dimensional persistence diagrams for three random seeds. b. Representatives of the most two most persistent features superimposed on a 2D PCA of the dataset. These always corresponded to the full shape and one of the two circle segments. For the first two random seeds, some points are distorted in such a way that they form a bridge in the 2D PCA, while in the third there is not such bridge and the second most persistent feature is much less persistent. Note that this is just a 2D PCA, in particular, much of the noise in 50D is not visible. A similar explanation applies for the diffusion distance in panel b of Figure 6.
Figure S18: $t$-SNE embeddings with perplexity $\rho=8$ and 1D persistence diagrams of the embedding for a circle in ambient $\mathbb{R}^{50}$ with Gaussian noise of low standard deviation $\sigma$. Perplexity $\rho=8$ is rather small, such that each embedding point only has attraction to very few other points. In the noiseless setting this very sparse attraction is only among immediate neighbors along the circle. This allows the embedding to have spurious curves. For higher noise, the sparse attraction pattern is less regular and local such that the spurious curves disappear. The more spurious curves the embedding has, the more high persistent features, given by bottlenecks in the curvy embedding, exist. This explains the dip for the $t$-SNE $\rho=8$ curve in panel g of Figure S19.
@@
Figure S22: Loop detection score for persistent homology with various distances on the noisy eyeglasses dataset in ambient $\mathbb{R}^{50}$. Only Fermat distance, embedding methods and spectral methods could handle low noise levels. Some embedding methods performed very well, but UMAP struggles in the low noise setting. The reason for the high random seed variability for effective resistance with $k=15$ is depicted in Figure S17.
@@
Figure S26: 2-loop detection score for persistent homology with various distances on a noisy torus in ambient $\mathbb{R}^{50}$. All methods struggled here, and only DTM, core, $t$-SNE graph and UMAP graph improved noticeably over the Euclidean distance. On a denser sampled torus effective resistance and diffusion distance outperformed other methods (Figure S27).

### v3

@@ Appendix G Effective resistance integrates diffusion distances
This is an application of the geometric series. Since $G$ is not bipartite, the largest eigenvalue $\mu_{n}$ of its Laplacian $L^{\text{sym}}$ is smaller than $2$ [17, Lemma 1.7]. Recall that by Eq. (4) the diffusion distance for $t$ diffusion steps is given by
@@
Corollary G.2.
The sum of the squared diffusion distances over all time points $t/2$ for $t=2,3,\dots$ equals the corrected effective resistance. Formally, if $G$ is a connected non-bipartite graph, we have
@@
For $x\in(-1,1)$, starting a geometric series at the $m$-th term yields $\sum_{t=m}^{\infty}x^{t}=\frac{x^{m}}{1-x}$. By Proposition F.2, we have
@@
By the same argument as in the proof of Proposition G.1, this shows the first part of the Corollary.
The statement for diffusion distance after $t=0$ steps is an easy computation from the original definition, Eq. (2):
@@
and the expression of the naive effective resistance as full geometric series of squared diffusion distances in Proposition G.1. ∎
Appendix H Diffusion pseudotime
@@
The difference is the $v_{1},\dots,v_{n}$’s. In the original publication [36], they were the normalized eigenvectors of the random walker graph Laplacian $D^{-\frac{1}{2}}L^{\text{sym}}D^{\frac{1}{2}}$. These are given by $v_{l}=D^{-\frac{1}{2}}u_{l}/\|D^{-\frac{1}{2}}u_{l}\|$. Wolf et al. [88] introduced a version using the eigenvectors of the symmetric graph Laplacian $L^{\text{sym}}$, so that $v_{l}=u_{l}$. Closest to the corrected resistance distance is the case where $v_{l}=D^{-\frac{1}{2}}u_{l}$ are non-normalized. We refer to these three versions as “rw”, “sym” and “symd”.
The only difference between the “symd” version of DPT and corrected effective resistance is that the former decays eigenvalues with $(1-\mu_{l})/\mu_{l}$, while the latter decays it with $(1-\mu_{l})/\sqrt{\mu_{l}}$, so that DPT decays large eigenvalues more strongly than corrected effective resistance (Figure 5).
@@
H.1 Performance of diffusion pseudotime and potential distance
Here we compare the spectral methods of our main benchmark to the three versions of DPT and potential distance.
For DPT, we used all three versions defined above and used $k\in\{15,100\}$ nearest neighbors.
@@ H.1 Performance of diffusion pseudotime and potential distance
Overall, all spectral methods performed very similarly on the 1D datasets circle, linked circles, and eyeglasses (Figures S7, S8), as well as on the single-cell datasets (Figure S9).
@@
Figure S9: Comparison of all spectral methods on the single cell datasets. They all achieve similar performance.
Appendix I Details on the distances used in our benchmark
@@
$\theta=\begin{cases}(\text{dtm}_{i}+\text{dtm}_{j}+d_{ij})/2&\text{ if }\xi=1\\ \left.\sqrt{\big{(}(\text{dtm}_{i}+\text{dtm}_{j})^{2}+d_{ij}^{2}\big{)}\cdot\big{(}(\text{dtm}_{i}-\text{dtm}_{j})^{2}+d_{ij}^{2}\big{)}}\middle/(2d_{ij})\right.&\text{ if }\xi=2\\ \max(\text{dtm}_{i},\text{dtm}_{j},d_{ij}/2)&\text{ if }\xi=\infty.\end{cases}$ (62)
@@
We computed the effective resistance on the $\text{s}k\text{NN}$ graph. Following the analogy with resistances in an electric circuit, if the $\text{s}k\text{NN}$ graph is disconnected, we computed the effective resistance separately in each connected component and set resistances between components to $\infty$. The uncorrected resistances were computed via the pseudoinverse of the unnormalized graph Laplacian [28]
@@
where $l_{ij}^{\dagger}$ is the $ij$-th entry of the pseudoinverse $L^{\dagger}$ of the unnormalized $\text{s}k\text{NN}$ graph Laplacian $L=D-A$. The pseudoinverse inverts all non-zero eigenvalues. Denote the eigenvalue decomposition of $L$ by $L=V\Lambda V^{T}$, where $V=(v_{1},\dots,v_{n})^{T}$ is the matrix of eigenvectors of $L$ and $\Lambda$ is the diagonal matrix of their eigenvalues $\lambda_{1},\dots,\lambda_{n}$. Then $L^{\dagger}=V\Lambda^{\dagger}V^{T}$, where $\Lambda^{\dagger}$ is the diagonal matrix with entry $1/\lambda_{i}$ if $\lambda_{i}>0$ and 00 otherwise. We can use this to derive a similar coordinate expression as in Eq. 5, but based on the unnormalized graph Laplacian ([28, 4.B]). Let $e_{i}$ be the $i$-th standard basis vector
$\tilde{d}^{\text{eff}}_{ij}$ $=l_{ii}^{\dagger}-2l_{ij}^{\dagger}+l_{jj}^{\dagger}$
$=(e_{i}-e_{j})^{T}L^{\dagger}(e_{i}-e_{j})$
$=(e_{i}-e_{j})^{T}V\Lambda^{\dagger}V^{T}(e_{i}-e_{j})$
$=\left(\sqrt{\Lambda^{\dagger}}V^{T}e_{i}-\sqrt{\Lambda^{\dagger}}V^{T}e_{j}\right)^{T}\left(\sqrt{\Lambda^{\dagger}}V^{T}e_{i}-\sqrt{\Lambda^{\dagger}}V^{T}e_{j}\right)$
$=\|\hat{e}_{i}^{\text{eff}}-\hat{e}_{j}^{\text{eff}}\|^{2}\text{ where}$
$\hat{e}_{i}^{\text{eff}}$ $=\left(\frac{v_{2,i}}{\sqrt{\lambda_{i}}},\dots,\frac{v_{n,i}}{\sqrt{\lambda_{n}}}\right).$ (68)
@@
For the weighted version of effective resistance, each edge in the $\text{s}k\text{NN}$ graph was weighted by the inverse of the Euclidean distance. We experimented with the weighted and unweighted versions, but only reported the unweighted version in the paper as the difference was always minor. We also experimented with the unweighted and uncorrected version and saw that correcting is crucial for high noise levels (Figure S20). We used $k~{}\in\{15,100\}$.
Both forms of the effective resistance can be written as squared distances between certain embedding points (5), (6). Nevertheless, the uncorrected effective resistance is a proper metric [32, Corollary 2.4.]. The corrected version in general is not a proper metric (Proposition F.3).
@@
It is clear from the above definition of the diffusion distance that it satisfies the triangle inequality, but it can fail to be positive on distinct points (Proposition F.3).
@@
Diffusion pseudotime [36] also integrates the different time scales of the diffusion distance, but on the level of the transition matrices, rather than on the level of the distances themselves (Proposition G.2). There are multiple variants of diffusion pseudotime. They are all computed as $d^{\textnormal{dpt}}_{ij}=\|e_{i}^{\textnormal{dpt}}-e_{j}^{\textnormal{dpt}}\|$ for
@@
We illustrate the distances used in the main benchmark in Figures S10 and S11.
Appendix J Datasets
J.1 Synthetic datasets
@@
J.2 Single-cell datasets
@@
Appendix K Hyperparameter selection
@@
The hyperparameters for Figures S10 and S11 are the same as those used in Figure 6. The hyperparameters for Figure S3 are given in Table S3 and those for Figure S16 in Table S4. All other supplementary figures either do not depend on hyperparameters or detail them directly in the figure or the caption.
@@
Dataset Fermat DTM Eff. res. Diffusion
Circle $p=2$ $k=4,p=2,\xi=1$ $k=100$ $k=15,t=8$
Eyeglasses $p=7$ $k=4,p=2,\xi=1$ $k=15$ $k=15,t=8$
Linked circles $p=5$ $k=4,p=2,\xi=1$ $k=100$ $k=15,t=8$
@@
Appendix L Hyperparameter sensitivity
While all distances other than the Euclidean distance have hyperparameters and we show results for the best hyperparameter setting in most figures, hyperparameter selection does not pose a serious problem for diffusion distance and effective resistance. We only tuned hyperparameters very mildly for these methods (4 settings for diffusion distances, only 2 for effective resistance, as opposed to 24 for the competing method DTM). Moreover, we conduced a more fine-grained sensitivity analysis and found the performance of diffusion distances and effective resistances to be robust to the exact value of their hyperparameters. For this sensitivity analysis we computed the area under the noise-level / detection score curves for the circle, the interlinked circles and the eyeglasses dataset in $50$ ambient dimensions with $k=4,15,30,45,60,75,90,105,120,135,150$ nearest neighbors and $t=2,4,8,16,32,64,128,256$ diffusion steps. For all but the most extreme hyperparameter values both effective resistance and diffusion distances strongly outperformed the Euclidean distance (Figure S15). We observed a trade-off between the optimal number of neighbors $k$ and diffusion steps $t$ for the diffusion distance, since both control how fast the diffusion can progress across the dataset.
@@
Appendix M Implementation details
We computed persistent homology using the ripser [4] project’s representative-cycles branch at commit 140670f to compute persistent homologies and representative cycles. We used coefficients in $\mathbb{Z}/2\mathbb{Z}$. To compute $k$NN graphs, we used the PyKeops package [13]. The rest of our implementation is in Python. Our code is available at https://github.com/berenslab/eff-ph/tree/neurips2024.
@@
The run time of persistent homology vastly dominated the time taken by the distance computation. The persistent homology run time depended most strongly on the sample size $n$, the dataset, and on the highest dimensionality of holes. The difference between distances was usually small. However, we observed that there were some outliers, depending on the noise level and the random seed, that had much longer run time. Overall, we found that methods that produce many pairwise distances of the same value (e.g., because of infinite distance in the graph affinities or maximum operations like for DTM with $p=\infty,\xi=\infty$) often had a much longer run time than other settings. We presume this was because equal distances led to many simplices being added to the complex at the same time. We give exemplary run times in Table S5.
@@
Appendix N Effect of outliers
Persistent homology with the Euclidean distance is known to be sensitive to outliers. Methods such as DTM were introduced to handle this issue. Here we show that spectral methods can also handle outliers well. Moreover, in high ambient dimensionality outliers are distributed over the large volume and hence are very sparse, making them less of a problem.
@@
In low dimensionality, Euclidean distance suffered in the low Gaussian noise setting already when only $50$ outliers were added. Adding $100$ or $200$ outliers severely lowered the detection score for Euclidean distance across the entire range of Gaussian noise strength. Fermat distances suffered from high random seed variability when adding outliers in low ambient dimension. Diffusion distance and effective resistance were much more outlier-resistant than the Euclidean distance and were only affected by $200$ outliers. Even then they performed better than the Euclidean distance without any outliers. DTM excelled in this setting, being completely insensitive to outliers and achieving top score for all noise levels (Figure S16a – d).
The volume of the bounding box in $d=50$ ambient dimensions is much larger and thus the same number of outlier points are distributed much more sparsely. In particular, it is much less likely that an outlier happens to fall into the middle of the circle. As a result, even Euclidean and Fermat distances were very outlier-robust in $d=50$ ambient dimensions (Figure S16e – g). Similarly, DTM’s performance did not change at all in the face of outliers. However, all three methods suffered strongly from the high-dimensional Gaussian noise.
As diffusion distance and effective resistance in our implementation rely on the unweighted $k$NN graph, they were somewhat more susceptible to outliers. Performance of diffusion distance decreased steadily with the number of outliers in high ambient dimension. When outliers were present, it performed worse than the Euclidean distance in the low Gaussian noise setting, but much better in the high Gaussian noise setting, even for $200$ outliers. Effective resistance performed best overall, deteriorating only slightly in the low Gaussian noise setting when outliers were added. Both spectral methods clearly outperformed other methods in the high-Gaussian noise regime even in the presence of numerous outliers.
@@
Figure S16: Loop detection performance of various methods on the noisy circle in the presence of outliers in low- and high-dimensional ambient space. Outliers were sampled uniformly from an axis-aligned cube around the data. a – d. In low ambient dimension ($d=2$) adding outliers hurt the performance of the Euclidean and Fermat distances, but barely affected the performance of the spectral methods and not at all DTM’s excellent performance. e – g. In high ambient dimensionality ($d=50$) outliers did not further decrease the weak performance of non-spectral methods. Diffusion distance was somewhat outlier-sensitive, but could still detect the loop structure in the high Gaussian noise setting. Effective resistance performed best overall and was not very outlier-sensitive.
Appendix O UMAP with higher embedding dimension
@@
Following the original publication [53], UMAP is typically used to embed data into two dimensions. This is an obvious issue for datasets sampled from manifolds which are not embeddable into two dimensions, such as the sphere or the torus. Therefore, we also experimented with the less common approach of higher embedding dimensionality ($3$, $5$, $10$) for UMAP. Save for the successful detection of the sphere’s void with at least three embedding dimensions, we observed few consistent effects of the embedding dimension either on the toy (Figure S17) or on the real data (Figure S19). Nevertheless, on the linked circles dataset, which is not embeddable into two dimensions, we saw a small improvement for UMAP when going beyond two embedding dimensions for the low noise setting.
@@
Independent of the embedding dimension, UMAP struggled in the low-noise setting on the eyeglasses dataset for $k=15$. Moreover, we observed very poor performance both for loop and void detection on the torus. We believe the reasons may be UMAP’s tendency to over-fragment the manifold [22] and UMAP’s use of a heavy-tailed kernel. We visualized three-dimensional UMAP embeddings of the noiseless eyeglasses, sphere, and torus for $k=15$ in Figure S18. UMAP left a gap in the embedding of the eyeglasses dataset, so that the true loop only had the second highest persistence. Short-cutting at the bottleneck yielded the most persistent loop. For the torus, the surface of the embedding was very fragmented, preventing the detection of any void. While the main loop of the torus was detected well, the second most persistent detected loop was already in the noise cloud as the fragmented surface of the embedding allowed for many fairly persistent loops. In a similar way, the surface of the sphere’s embedding got fragmented, leading to many loops. Our score does not penalize this, because the detection score for $m=1$ loop is low as there is nearly no gap in persistence between the most and the second-most persistent loop. Higher levels of noise and also higher $k$ could overcome UMAP’s over-fragmentation tendency for the eyeglasses dataset. This may be due to high-dimensional noise better matching UMAP’s heavy-tailed kernel. However, it did not resolve the over-fragmentation for the torus.
@@
The runtime of UMAP scales linearly with the embedding dimension, but typical $t$-SNE implementations scale exponentially. In fact, the implementation we used here, openTSNE [63], does not implement embedding dimensions higher than two, which is why we explored higher-dimensional embeddings only for UMAP.
@@
Figure S19: Detection scores for UMAP with different embedding dimensions on the single-cell datasets with some other methods for reference. Changing the embedding dimension did not have a consistent effect on the detection scores, while higher embedding dimension usually hurt for Laplacian Eigenmaps.
Appendix P Additional figures
Figure S20: Loop detection score on noisy $S^{1}\subset\mathbb{R}^{50}$ for various versions of effective resistance. There was little difference between using the weighted $k$NN graph, unweighted $k$NN graph, and using the square root of effective resistance based on the unweighted $k$NN graph. The latter got filtered out for high noise levels. Using $k=100$ instead of $k=15$ helped only marginally in this dataset. The uncorrected (naive) version of effective resistance collapsed already at very small noise levels.
Figure S21: Illustration for the random seed variability of effective resistance with $k=15$ on the noisy eyeglasses dataset in $\mathbb{R}^{50}$ with $\sigma=0.075$. This refers to Figure 7b. a. One-dimensional persistence diagrams for three random seeds. b. Representatives of the most two most persistent features superimposed on a 2D PCA of the dataset. These always corresponded to the full shape and one of the two circle segments. For the first two random seeds, some points are distorted in such a way that they form a bridge in the 2D PCA, while in the third there is not such bridge and the second most persistent feature is much less persistent. Note that this is just a 2D PCA, in particular, much of the noise in 50D is not visible. A similar explanation applies for the diffusion distance in Figure 7b.
Figure S22: $t$-SNE embeddings with perplexity $\rho=8$ and 1D persistence diagrams of the embedding for a circle in ambient $\mathbb{R}^{50}$ with Gaussian noise of low standard deviation $\sigma$. Perplexity $\rho=8$ is rather small, such that each embedding point only feels attraction to very few other points. In the noiseless setting this very sparse attraction is only among immediate neighbors along the circle. This makes the embedding to have spurious curves. For higher noise, the sparse attraction pattern is less regular and less local such that the spurious curves disappear. The more spurious curves the embedding has, the more high persistent features, given by bottlenecks in the curvy embedding, exist. This explains the dip for the $t$-SNE $\rho=8$ curve in Figure S23g.
@@
Figure S26: Loop detection score for persistent homology with various distances on the noisy eyeglasses dataset in ambient $\mathbb{R}^{50}$. Only Fermat distance, geodesics, $t$-SNE and UMAP, and spectral methods outperformed the Euclidean distance, but UMAP struggled in the low noise setting. The reason for the high random seed variability for effective resistance with $k=15$ is depicted in Figure S21.
@@
Figure S30: 2-loop detection score for persistent homology with various distances on a noisy torus in ambient $\mathbb{R}^{50}$. All methods struggled here, and only DTM, core, $t$-SNE graph and UMAP graph improved noticeably over the Euclidean distance. On a denser sampled torus effective resistance and diffusion distance outperformed other methods (Figure S31). Using fewer diffusion steps improved the performance on the torus (Figure S32).